# Spurious reconstruction from brain activity

### v1

@@ Abstract
The rapid advances in brain decoding, particularly visual image reconstruction, have sparked discussions about the potential societal implications and ethical considerations surrounding neurotechnology. As these methods aim to recover perceived images from brain activity and achieve prediction over diverse images beyond training samples (zero-shot prediction), it is crucial to critically assess their capabilities and limitations to prevent misguided public expectations and inform future regulations. Our case study of recent text-guided reconstruction methods, which leverage a large-scale dataset (the Natural Scene Dataset, NSD) and text-to-image diffusion models, reveals significant limitations in their generalizability. We found a notable decrease in performance when applying these methods to a different dataset, which was designed to prevent category overlaps between training and test sets. UMAP visualization of the text features with NSD images showed a limited diversity of distinct semantic and visual clusters, with substantial overlap between training and test sets. Formal analysis and simulations demonstrated that clustered training samples can lead to “output dimension collapse,” restricting the output feature dimensions predictable from brain activity. Diversifying the training set to ensure a broader feature distribution improved generalizability beyond the trained clusters. However, text features alone are insufficient for a complete mapping to the visual space, even if perfectly predicted from brain activity. We argue that recent photo-like reconstructions may primarily be a blend of classification into trained categories and the generation of convincing yet inauthentic images through text-to-image diffusion (hallucination). To achieve genuine zero-shot prediction, diverse datasets and compositional representations spanning the image space are essential. As neurotechnology advances, engaging in interdisciplinary discussions involving neuroscientists, ethicists, policymakers, and the public is crucial to ensure responsible development and application of these techniques. These discussions should be grounded in a clear understanding of the current capabilities and limitations of the technology, as well as a careful consideration of the potential ethical and societal impacts.
Keywords Visual image reconstruction  $\cdot$ brain decoding  $\cdot$ zero-shot prediction  $\cdot$ methodology  $\cdot$ generative AIs
1 Introduction
@@
Brain decoding has been widely used in the neuroscience field, revealing specific contents of the mind (Haxby et al., 2001; Kamitani and Tong, 2005; Soon et al., 2008; Horikawa et al., 2013). As brain decoding is sometimes referred to as “mind-reading” in popular media (Somers, 2021; Whang, 2023; Raasch, 2023), it has attracted significant attention beyond the scientific community due to its potential for real-world applications in medicine and industry. Such neurotechnology has also started to affect future ethical discussions and legal regulations (UNESCO, 2023). To prevent misguiding public expectations and policies, scientists need to carefully assess the current status of brain decoding techniques and clarify the possibilities and limitations.
@@
One of the major challenges in brain decoding is the limited amount of brain data we can collect. The current brain measurement devices are costly, yielding far less brain data than the amounts typically used in image or text processing within the field of computer science and AI (Deng et al., 2009; Schuhmann et al., 2022). Although we have gradually increased the amount of brain data per subject (Van Essen et al., 2012; Allen et al., 2022; Naselaris et al., 2021; Hebart et al., 2023; Xu et al., 2024), it remains impractical to collect brain data covering the full range of cognitive states and perceptual experiences. Consequently, classification-based decoding approaches, primarily developed in the early stages of this field, are insufficient to uncover the neural code in general or natural conditions since decodable information is confined to the same stimuli or predefined categories used in a training phase.
@@
To overcome this limitation, several decoding methods have been proposed to enable the prediction of novel contents from brain activities that are not encountered during the training phase. Kay et al. (2008) proposed a general visual decoding approach via a statistical encoding model that predicted fMRI voxel values from image features. It successfully identified novel test images from a set of 1000 candidates. Mitchell et al. (2008) utilized co-occurrence rates of specific verb sets for nouns to predict fMRI brain activity during the perception of line drawing images, demonstrating the ability to predict unseen nouns. Brouwer and Heeger (2009) constructed a color-tuning model and predicted brain activity while the subjects were presented with color stimuli. As their training stimuli cover most of the color space, their methods can successfully identify novel colors not included in the training dataset. Horikawa and Kamitani (2017) utilized deep neural network (DNN) features to decode brain activity measured while subjects perceived natural images. They showed successful prediction of novel object categories not encountered during the training phase.
@@
In the field of machine learning, “zero-shot” prediction refers to the ability of a model to accurately predict or classify novel contents that were not encountered during the training phase (Larochelle et al., 2008; Palatucci et al., 2009). This technique has found widespread application across various domains, including image classification (Radford et al., 2021), image generation (Ramesh et al., 2021), and natural language processing (Brown et al., 2020). The concept of zero-shot prediction can be considered analogous to brain decoding techniques that aim to interpret brain activity patterns associated with previously unseen stimuli or experiences. Both approaches seek to generalize knowledge gained from a limited set of training data to novel situations, enabling the interpretation of new information without explicit prior exposure. To achieve effective zero-shot prediction, the model often utilizes a compositional representation of the output (Lake et al., 2017; Higgins et al., 2018). Compositional representation refers to the ability to understand and generate novel combinations of previously learned concepts or features. By learning the underlying structure and relationships between different elements, the model can generalize its knowledge to new, unseen instances.
@@
Visual image reconstruction is another prominent example of zero-shot prediction in brain decoding. This task aims to recover perceived novel images that were not encountered during the training phase, effectively reconstructing visual experiences from brain activity patterns (Stanley et al., 1999; Miyawaki et al., 2008). As our perceptual visual experiences cannot be fully covered by limited brain data, reconstruction methods require strong generalizability. Miyawaki et al. (2008) conducted a study that demonstrated the reconstruction of perceived arbitrary 10x10 binary-contrast images from brain activity. They built multiple modular decoders to predict the local contrasts of each location and combined their predictions. This approach leverages the compositional representation of the visual field, which is organized retinotopically in the early visual cortex. Incorporating cortical organization into the model’s architecture can significantly improve its ability to perform zero-shot prediction and reconstruct novel visual experiences from brain activity. Although the training stimuli were only 400 random images, it was possible to reconstruct an arbitrary image from a set of possible $2^{100}$ instances, including geometric shapes such as crosses and alphabets. Similarly, Shen et al. (2019b) replaced local decoders with DNN feature decoders. Although their training stimuli were 1200 natural images, they demonstrated reconstructing novel images, including artificial images, which were not part of the training set. These successes suggest that the proposed reconstruction models capture rich and comprehensive information about the general aspects of the neural code, beyond merely the information defined by the training data (Kriegeskorte and Douglas, 2019). Developing reliable reconstruction methods also enables further analysis of subjective visual experiences, such as visual imagery (Shen et al., 2019b), attention (Horikawa and Kamitani, 2022), and illusion (Cheng et al., 2023). Enabling the decoding of novel brain states that were never encountered during the training phase can be a promising approach to neural mind-reading (Kamitani and Tong, 2005).
Visual image reconstruction pipelines can be divided into three main components: translator, latent features, and generator (Figure 1). The translator serves to convert a brain activity pattern into a latent feature space via linear regression (Shen et al., 2019b; Seeliger et al., 2018; Mozafari et al., 2020; Ozcelik et al., 2022; Kavasidis et al., 2017), or nonlinear transformation (Qiao et al., 2020). Latent features serve as surrogate representations of perceived visual images. While the local contrasts in Miyawaki et al. (2008) can be seen as a primitive form of latent features, recent studies often use DNN features, such as the intermediate output of recognition models (Shen et al., 2019b; Cheng et al., 2023). The generator visualizes translated latent features into images. Some studies have used the pretraining image generative models for the generator module (Mozafari et al., 2020; Qiao et al., 2020; Ozcelik et al., 2022; Kavasidis et al., 2017). Image optimization can also be regarded as a generator (Shen et al., 2019b). Direct mapping from brain activity to images using DNNs can also be considered to implicitly contain these components (Shen et al., 2019a; Beliy et al., 2019; Ren et al., 2021; Gaziv et al., 2022; Lin et al., 2022; Chen et al., 2023a).
Figure 1: An overview of the visual image reconstruction pipeline. Brain activity recorded while perceiving a visual image is fed into a translator module. The translator maps the brain activity to latent features. The generator module converts the latent features into an image corresponding to the original visual stimulus.
@@
Recent advances in generative AI, particularly in text-to-image generation, have naturally given rise to expectations that these techniques could provide a shortcut for visual image reconstruction by leveraging semantic representations. In addition, there has been a growing trend towards collecting neural datasets using a wide range of diverse visual and semantic content. This shift aims to capture a more comprehensive and ecologically valid representation of the human experience (Naselaris et al., 2021). Researchers have started to collect large-scale datasets, such as the Natural Scene Dataset (NSD) (Allen et al., 2022) and the THINGS-fMRI dataset (Hebart et al., 2023), which incorporate a broader range of fMRI data induced by diverse visual stimuli with text or category annotations. Notably, recent studies leveraging the latest generative AI techniques with semantic representation and large-scale datasets have reported photo-like reconstructions from brain activity (Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Scotti et al., 2023; Bai et al., 2023; Chen et al., 2023b; Lan et al., 2023; Benchetrit et al., 2024). They utilized contrastive language-image pretraining (CLIP) features (Radford et al., 2021) as latent features and text-to-image diffusion models (Ramesh et al., 2021; Xu et al., 2022) as generators.
@@
While these recent approaches show promising results, it remains uncertain whether these methods truly achieve zero-shot reconstruction due to several factors. The complex model architectures employed in these studies, along with the use of a large-scale dataset, make it challenging to interpret and understand the underlying mechanisms driving the reconstruction process. To fully assess the zero-shot prediction capabilities of these approaches, it is essential to rigorously test their generalizability across different datasets and to provide detailed analyses of the individual model components. This includes evaluating the performance of the translators, latent features, and generators used in these methods. Furthermore, the characterization of the diversity of stimuli in the datasets and the latent representations has not been thoroughly explored. It is unclear whether the recently proposed datasets, such as the NSD, are optimally designed to capture the full range of human visual experiences and to support the development of truly generalizable prediction models.
@@
In the following, we begin with a case study that critically tests the methods proposed by Takagi and Nishimoto (2023b) and Ozcelik and VanRullen (2023), which originally used the NSD. Our findings reveal several issues, including the failure of replication with another dataset specifically designed to avoid overlap between training and test sets (Shen et al., 2019b), and the questionable post-hoc selection procedures of Takagi and Nishimoto (2023b) that can produce convincing reconstructions even with random brain data. We identify the lack of diversity represented by the limited number of clusters in the NSD dataset as a potential factor contributing to these issues. We also demonstrate the failure of zero-shot prediction in the feature space and the inability to recover a stimulus from its latent features, suggesting a misspecification of the model. Based on these findings, we conclude that the apparent photo-like reconstructions are essentially a result of classification to clusters shared between training and test sets, combined with hallucinations by the generative model.
@@
In the formal analysis and simulation section, we aim to uncover the general factors underlying the issues and observations highlighted in the case study. We first describe the phenomenon of “output dimension collapse” in the translation from brain activity to the latent feature space. We show that the regression model trained on the clustered outputs becomes specialized to the training examples and its output collapses into a subspace of the training dataset. Our simulations with clustered data demonstrate that out-of-sample prediction can be achieved with an increasing number of training clusters, suggesting that zero-shot prediction is possible if stimulus diversity and a compositional representation are available. We also explore the preservation of image information at hierarchical layers of DNNs and discuss the caveats associated with evaluating reconstructions using identification metrics alone. Finally, we provide speculation on how we are fooled by the seemingly realistic reconstructions generated by AI models.
@@
2 Results
2.1 Case study
@@
We investigated two types of recent generative AI-based reconstruction methods, StableDiffusionReconstruction (Takagi and Nishimoto, 2023b) and Brain-diffuser (Ozcelik and VanRullen, 2023), as well as their validation dataset, the NSD (Allen et al., 2022). Both reconstruction methods utilize CLIP features (Radford et al., 2021) to effectively apply recent text-to-image diffusion models in visual image reconstruction analysis. CLIP text features are obtained from the average of five text annotations corresponding to the stimulus image. These text annotation information are only used during training to map the brain activity into CLIP text features. In the test phase, they directly predict CLIP text features from brain activity during image perception. Henceforth, these two reconstruction methods will be referred to together as text-guided reconstruction methods.
@@
The StableDiffusionReconstruction method (Takagi and Nishimoto, 2023b) uses the components of the Stable Diffusion model (Ramesh et al., 2021), the VAE features (Kingma and Welling, 2014) and CLIP text features as latent features. Similarly, the Brain-diffuser method (Ozcelik and VanRullen, 2023) utilizes the component of another type of diffusion model (Xu et al., 2022), CLIP text features, and CLIP vision features as well as VDVAE features (Child, 2021) as latent features. Both methods first generate initial images from translated VAE/VDVAE features. These initial images are then passed through the image-to-image pipeline of the diffusion model conditioned on the translated CLIP features, producing the final reconstructed images. They validated the reconstruction performance using the NSD dataset, preparing training and test data based on the data split provided by the NSD study. Thanks to the authors of these studies who made efforts to make their datasets and scripts publicly available, we were able to conduct our replication analysis effectively. We compared the reconstructed results of these two text-guided reconstruction methods with those from a previous image reconstruction method, iCNN (Shen et al., 2019b). For more information on datasets and reconstruction methods, see Material and Methods (“Datasets” or “Reconstruction methods”) or the original studies (Allen et al., 2022; Shen et al., 2019b; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023).
@@
2.1.1 Observations: Failed replication and convincing reconstruction from random data
@@
We first confirmed the reproduction of the findings of the text-guided reconstruction methods, particularly those of the Brain-diffuser study (Ozcelik and VanRullen, 2023) when applied to the NSD dataset (Figure 2a). The reconstructed images generated by the Brain-diffuser method effectively captured most of the layout and semantics of the test images. Although the reconstructed images produced by the StableDiffusionReconstruction method (Takagi and Nishimoto, 2023b) showed slightly degraded performance compared to the original paper, they still successfully captured the semantics of the test images. Notably, the iCNN method (Shen et al., 2019b) also performed well on the NSD dataset, with reconstructed images capturing the dominant structures of the objects within the images, consistent with the findings reported in the original study.
To further investigate the generalizability of the text-guided reconstruction methods, we attempted to replicate their performance using a different dataset, Deeprecon, which was originally collected for the study by Shen et al. (2019b). The Deeprecon dataset was specifically designed to avoid overlap between training and test sets, making it an ideal benchmark for evaluating the zero-shot prediction capabilities of reconstruction methods. However, the original Deeprecon dataset lacked the text annotations required by the text-guided reconstruction methods. To facilitate a fair comparison, we collected five text annotations for each training stimulus in the Deeprecon dataset through crowd-sourcing and used them to generate CLIP text features.
Despite the inclusion of these text annotations, the text-guided reconstruction methods failed to achieve the same level of performance on the Deeprecon dataset as they did on the NSD dataset. The reconstructed images produced by the text-guided methods exhibited photo-like appearances but suffered from largely degraded quality compared to their performance on the NSD dataset (Figure 2b). Notably, the text-guided reconstruction methods generated photo-like reconstructions even for simple geometric shapes present in the Deeprecon dataset, which deviated significantly from the original stimuli. In contrast, the iCNN method (Shen et al., 2019a) consistently provided faithful reconstructions for both the NSD and Deeprecon datasets, despite its simplicity compared to the text-guided methods.
@@
Figure 2: Image reconstruction results. (a) Reconstruction results for the NSD. The first row shows the original test images, followed by reconstructions using StableDiffusionReconstruction, Brain-Diffuser, and iCNN methods. (b) Reconstruction results for the Deeprecon dataset. The format is the same as in (a).
@@
Upon further investigation, we noted a questionable post-hoc image selection procedure. In the StableDiffusionReconstruction study (Takagi and Nishimoto, 2023b), they presented the reconstruction results by the following procedure: “We generated five images for each test image and selected the generated images with highest PSM” (PSM means perceptual similarity metrics), as illustrated in Figure 3a. This selection might lead readers or peer reviewers, particularly those not specialized in the brain decoding field, to overestimate the effectiveness of the methods and potentially lead to a distorted understanding of the actual reconstruction performance. Note that Brain-diffuser study (Ozcelik and VanRullen, 2023) did not executed such procedures and in their following report (Takagi and Nishimoto, 2023a), they updated the image presentation procedure more fairly as: “we generated five images with different stochastic noise and selected three images randomly.”
@@
Figure 3: Questionable post-hoc selection. (a) Procedure in the StableDiffusionReconstruction method. Five images were generated, and the one that most closely resembled the test images was selected. (b) Reconstruction using shuffled brain activity. The left panels show the images generated from the VAE features given shuffled brain activity and CLIP text features given original brain activity. The right panels show the images generated from the VAE and CLIP text features given shuffled brain activity.
@@
To examine the impact of this post-hoc selection procedure, we conducted an experiment using random brain data (Figure 3b). Instead of feeding the test brain data into pretrained translators, we shuffled the activities within each voxel of the NSD test set independently to create random brain data. Surprisingly, when the random brain data were input into the VAE feature translator, which contributes to producing initial images, plausible images were obtained by generating five images and selecting the best one. Even more strikingly, when the random brain data were input into both VAE and CLIP text feature translators, we still obtained convincing results by simply generating images five times and conducting the aforementioned selection. These observations are inexplicable since the artificially created brain data should completely lack any information related to the original visual stimuli.

### v3

@@ Abstract
Advances in brain decoding, particularly visual image reconstruction, have sparked discussions about the societal implications and ethical considerations of neurotechnology. As these methods aim to recover visual experiences from brain activity and achieve prediction beyond training samples (zero-shot prediction), it is crucial to assess their capabilities and limitations to inform public expectations and regulations. Our case study of recent text-guided reconstruction methods, which leverage a large-scale dataset (Natural Scene Dataset, NSD) and text-to-image diffusion models, reveals limitations in their generalizability. We found poor performance when applying these methods to a different dataset designed to prevent category overlaps between training and test sets. UMAP visualization of the text features with NSD images showed a limited diversity of semantic and visual clusters, with overlap between training and test sets. Formal analysis and simulations demonstrated that clustered training samples can lead to “output dimension collapse,” restricting predictable output feature dimensions. Simulations further showed that diversifying the training set improved generalizability. However, text features alone are insufficient for mapping to the visual space. We argue that recent realistic reconstructions may primarily be a blend of classification into trained categories and generation of inauthentic images through text-to-image diffusion (hallucination). Diverse datasets and compositional representations spanning the image space are essential for genuine zero-shot prediction. Interdisciplinary discussions grounded in understanding the current capabilities and limitations, as well as ethical considerations, of the technology are crucial for its responsible development.
Author summary
Brain decoding has gained attention not only in neuroscience but also in fields like computer science, and even policy-making. However, claims about its predictive performance often tend to be overly optimistic. Here, we critically examine recent brain decoding methods, particularly visual image reconstruction methods that use generative AIs to generate seemingly realistic images from brain activity. Our analyses reveal that the apparent success of these methods largely stems from exploiting similarities between training and test stimuli in the experiment and from the hallucination of recent generative AIs, failing to achieve the genuine zero-shot reconstruction of novel images. We introduce the concept of “output dimension collapse,” a fundamental limitation in predicting within a target space, and discuss conditions necessary for true zero-shot prediction in terms of compositional representations and image recovery from latent features. We also highlight the limitations of current evaluation metrics and provide a framework for interpreting seemingly realistic reconstructions. By examining current limitations and proposing improvement strategies, we aim to encourage rigorous development and evaluation of brain decoding techniques.
Introduction
@@
Brain decoding has been widely used in the neuroscience field, revealing specific contents of the mind [1, 2, 3, 4]. As brain decoding is sometimes referred to as “mind-reading” in popular media [5, 6, 7], it has attracted significant attention beyond the scientific community due to its potential for real-world applications in medicine and industry. Such neurotechnology has also started to affect future ethical discussions and legal regulations [8]. To prevent misleading public expectations and policies, scientists need to carefully assess the current status of brain decoding techniques and clarify the possibilities and limitations.
@@
One of the major challenges in brain decoding is the limited amount of brain data we can collect. The current brain measurement devices are costly, yielding far less brain data than the amounts typically used in image or text processing within the field of computer science and AI [9, 10]. Although we have gradually increased the amount of brain data per subject [11, 12, 13, 14, 15], it remains impractical to collect brain data covering the full range of cognitive states and perceptual experiences. Consequently, classification-based decoding approaches, primarily developed in the early stages of this field, are insufficient to uncover the neural code in general or natural conditions since decodable information is confined to the same stimuli or predefined categories used in a training phase.
@@
To overcome this limitation, several decoding methods have been developed to enable the prediction of novel contents from brain activities that are not encountered during the training phase. Kay et al. [16] proposed a general visual decoding approach via a statistical encoding model that predicted fMRI voxel values from image features. It successfully identified novel test images from a set of 1,000 candidates. Mitchell et al. [17] utilized co-occurrence rates of specific verb sets for nouns and built a computational model to predict fMRI voxel values while thinking about nouns in presented line drawing images. Their model demonstrated the ability to predict voxel values for novel nouns not seen during the training phase. Brouwer and Heeger [18] constructed a color-tuning model and predicted brain activity while the subjects were presented with color stimuli. As their training stimuli covered most of the color space, their methods successfully identified novel colors not included in the training dataset. Horikawa and Kamitani [19] utilized deep neural network (DNN) features to decode brain activity measured while subjects perceived natural images. They showed successful prediction of novel object categories not encountered during the training phase.
@@
In the field of machine learning, “zero-shot” prediction refers to the ability of a model to accurately predict or classify novel contents not encountered during the training phase [20, 21]. This capability has emerged in various applications across different domains, including image classification [22], image generation [23], and natural language processing [24]. The concept of zero-shot prediction can be considered analogous to brain decoding techniques that aim to interpret brain activity patterns associated with previously unseen stimuli or experiences. Both approaches seek to generalize knowledge gained from a limited set of training data to novel situations, enabling the interpretation of new information without explicit prior exposure. To achieve effective zero-shot prediction, the model often utilizes a compositional representation of the output [25, 26]. Compositional representation refers to the ability to understand and generate novel combinations of previously learned concepts or features. By learning the underlying structure and relationships between different elements, the model can generalize its knowledge to new, unseen instances.
@@
Visual image reconstruction is another prominent example of zero-shot prediction in brain decoding. This task aims to recover perceived novel images that were not encountered during the training phase, effectively reconstructing visual experiences from brain activity patterns [27, 28]. As our perceptual visual experiences cannot be fully covered by limited brain data, reconstruction methods require strong generalizability. Miyawaki et al. [28] conducted a study demonstrating the reconstruction of perceived arbitrary $10\times 10$ binary-contrast images from brain activity. They built multiple modular decoders to predict the local contrasts of each location and combined their predictions. This approach leverages the compositional representation of the visual field, which is organized retinotopically in the early visual cortex. Incorporating cortical organization into the model’s architecture can significantly improve its ability to perform zero-shot prediction and reconstruct novel visual experiences from brain activity. Although the training stimuli were only $400$ random images, it was possible to reconstruct an arbitrary image from a set of possible $2^{100}$ instances, including geometric shapes such as crosses and alphabets. Similarly, Shen et al. [29] replaced local decoders with DNN feature decoders. Although their training stimuli were $1,200$ natural images, they demonstrated reconstructing novel images, including artificial images, which were not part of the training set. These successes suggest that the proposed reconstruction models capture rich and comprehensive information about the general aspects of the neural code, beyond merely the information defined by the training data [30]. Developing reliable reconstruction methods also enables further analysis of subjective visual experiences, such as visual imagery [29], attention [31], and illusion [32]. Enabling the decoding of novel brain states that were never encountered during the training phase can be a promising approach to neural mind-reading [2].
Visual image reconstruction pipelines typically comprise three main components: translator, latent features, and generator (Fig 1). The translator converts brain activity patterns into a latent feature space, employing either linear regression [29, 33, 34, 35] or nonlinear transformation [36]. Latent features serve as surrogate representations of perceived visual images, evolving from primitive forms like local contrasts [28] to more sophisticated DNN features, such as intermediate outputs of recognition models [19, 29]. In our current work, we reframe this process as “translation” rather than “feature decoding,” a term we used in previous studies. This terminology acknowledges two important points: first, brain activity itself can be considered a latent representation of an image or perception of it, and second, it helps avoid potential ambiguity between image encoding/decoding and brain encoding/decoding processes. This new perspective conceptualizes the process as a translation between two latent spaces: from neural representation to machine representation.
Fig 1: Visual image reconstruction pipeline. The first step involves translating brain activity patterns into machine/AI latent representations. While this process typically involves brain decoding of latent features using machine learning, here we term it “translation.” This terminology acknowledges that brain activity itself can be viewed as a latent representation of an image, framing the process as a translation from neural to machine latent representations. In the second step, a generator module takes these translated latent features as input and converts them into a visual image that corresponds to the content represented by the original brain activity.
The generator visualizes translated latent features into images. Some studies have used pretraining image generative models for the generator module [34, 36, 35]. Image optimization can also be regarded as a generator [29]. End-to-end mapping from brain activity to images using DNNs can also be considered to contain these components implicitly or as a generator-only method [37, 38, 39, 40, 41, 42, 43].
@@
Recent advances in generative AI, particularly in text-to-image generation, have naturally given rise to expectations that these techniques could provide a valuable tool for visual image reconstruction by leveraging semantic representations. In addition, there has been a growing trend towards collecting neural datasets using a wide range of diverse visual and semantic content. This shift aims to capture a more comprehensive and ecologically valid representation of the human experience [13]. Researchers have started to collect large-scale fMRI datasets, such as the Natural Scene Dataset (NSD) [12] and the THINGS-fMRI dataset [14], which incorporate a broader range of brain data induced by diverse visual stimuli with text or category annotations. Notably, recent studies leveraging the latest generative AI techniques with semantic representation and large-scale datasets have reported realistic reconstructions from brain activity [44, 45, 46, 47, 48]. They utilized contrastive language-image pretraining (CLIP) features [22] as latent features and text-to-image diffusion models [23, 49] as generators.
@@
While these recent approaches show promising results, it remains uncertain whether these methods truly achieve zero-shot reconstruction due to several factors. The complex model architectures employed in these studies, along with the use of a large-scale dataset, make it challenging to interpret and understand the underlying mechanisms driving the reconstruction process. To fully assess the zero-shot prediction capabilities of these approaches, it is essential to rigorously test their generalizability across different datasets and to provide detailed analyses of the individual model components. This test includes evaluating the performance of the translators, latent features, and generators used in these methods. Furthermore, the characterization of the diversity of stimuli in the datasets and the latent representations has not been thoroughly explored. It is unclear whether the recently proposed datasets, such as the NSD, are optimally designed to capture the full range of human visual experiences and to support the development of truly generalizable prediction models.
@@
In the following, we begin with a case study that critically tests the methods proposed by Takagi and Nishimoto [44] and Ozcelik and VanRullen [45], which originally used the NSD. Our findings reveal several issues, including the failure to replicate the same level of reconstruction with another dataset specifically designed to avoid overlap between training and test sets [29], and the questionable post-hoc selection procedures of Takagi and Nishimoto that can produce convincing reconstructions even with random brain data. We identify the lack of diversity represented by the limited number of semantic clusters in the NSD dataset as a potential factor contributing to these issues. We also demonstrate the failure of zero-shot prediction in the latent feature space and the inability to recover a stimulus from its latent features, suggesting a misspecification of the model. Based on these findings, we conclude that the apparent realistic reconstructions are essentially a result of classification to clusters shared between training and test sets, combined with hallucinations by the generative model.
@@
In the formal analysis and simulation section, we aim to uncover the general factors underlying the issues and observations highlighted in the case study. We first describe the phenomenon of “output dimension collapse” in the translation from brain activity to the latent feature space. We show that the regression model trained on the clustered targets becomes specialized to the training examples, and its outputs collapse into a subspace of the training set. Our simulations with clustered data demonstrate that out-of-sample prediction can be achieved with an increasing number of training clusters, suggesting that zero-shot prediction is possible if stimulus diversity and a compositional representation are available. We also discuss the caveats associated with evaluating reconstructions using identification metrics alone and explore the preservation of image information at hierarchical layers of DNNs. Finally, we provide speculation on how we are fooled by the seemingly realistic reconstructions generated by AI models.
@@
Results
Case study
@@ Results
We investigated two types of recent generative AI-based reconstruction methods, StableDiffusionReconstruction [44] and Brain-Diffuser [45], as well as their validation dataset, the Natural Scene Dataset (NSD) [12]. Both reconstruction methods utilize CLIP features [22] to effectively apply recent text-to-image diffusion models in visual image reconstruction analysis. CLIP text features are obtained from the average of five text annotations corresponding to the stimulus image. This text annotation information is only used during training to map the brain activity into the CLIP text features. In the test phase, they directly translate text features from brain activity during image perception. Hence, these two reconstruction methods will be referred to together as text-guided reconstruction methods.
@@
The StableDiffusionReconstruction method [44] uses the components of the Stable Diffusion model [50], the VAE features [51], and CLIP text features as latent features. Similarly, the Brain-Diffuser method [45] utilizes the component of another type of diffusion model [49], CLIP text features, CLIP vision features, and VDVAE features [52] as latent features. Both methods first generate initial images from translated VAE/VDVAE features. These initial images are then passed through the image-to-image pipeline of the diffusion model conditioned on the translated CLIP features, producing the final reconstructed images. They validated the reconstruction performance using the NSD dataset, preparing training and test data based on the data split provided by the NSD study. Thanks to the authors of these studies who made efforts to make their datasets and scripts publicly available, we were able to conduct our replication analysis effectively. We compared the reconstructed results of these two text-guided reconstruction methods with those from a previous image reconstruction method, iCNN [29]. For more information on datasets and reconstruction methods, see Material and Methods (“Datasets” or “Reconstruction methods”) or the original studies [12, 29, 44, 45].
@@
Observations: Failed replication and convincing reconstruction from random data
@@
We first confirmed the reproduction of the findings of the text-guided reconstruction methods, particularly those of the Brain-Diffuser study [45] when applied to the NSD dataset (Fig 2A). The reconstructed images generated by the Brain-Diffuser method effectively captured most of the layout and semantics of the test images. Although the reconstructed images produced by the StableDiffusionReconstruction method [44] showed slightly degraded performance compared to the original paper, they still successfully captured the semantics of the test images. Notably, the iCNN method [29] also performed well on the NSD dataset, with reconstructed images capturing the dominant structures of the objects within the images, consistent with the findings reported in the original study.
@@
Fig 2: Comparison of image reconstruction results across datasets and methods. (A) Reconstruction results for the Natural Scene Dataset (NSD). The first row shows the original test images, followed by reconstructions using StableDiffusionReconstruction, Brain-Diffuser, and iCNN methods. (B) Reconstruction results for the Deeprecon dataset, presented in the same format as (A).
To further investigate the generalizability of the text-guided reconstruction methods, we attempted to replicate their performance using a different dataset, Deeprecon, which was originally collected for the study by Shen et al. [29]. The Deeprecon dataset was explicitly designed to avoid overlap between training and test sets, making it an ideal benchmark for evaluating the zero-shot prediction capabilities of reconstruction methods. However, the original Deeprecon dataset lacked the text annotations required by the text-guided reconstruction methods. To facilitate a fair comparison, we collected five text annotations for each training stimulus in the Deeprecon dataset through crowd-sourcing and used them to generate CLIP text features.
Despite including these text annotations, the text-guided reconstruction methods failed to achieve the same level of performance on the Deeprecon dataset as they did on the NSD dataset. The reconstructed images produced by the text-guided methods exhibited realistic appearances but suffered from largely degraded quality compared to their performance on the NSD dataset (Fig 2B). Notably, the text-guided reconstruction methods generated realistic reconstructions even for simple geometric shapes in the Deeprecon dataset, which deviated significantly from the original stimuli. In contrast, the iCNN method consistently provided faithful reconstructions for both the NSD and Deeprecon datasets, despite its simplicity compared to the text-guided methods. These results suggest that the text-guided methods may not generalize well across different datasets. The ability to generate realistic but inaccurate reconstructions, especially for simple shapes, indicates that these methods might rely more on training stimuli than on actual brain activity information.
@@
Upon further investigation, we noted a questionable post-hoc image selection procedure. In the StableDiffusionReconstruction study [44], they presented the reconstruction results by the following procedure: “We generated five images for each test image and selected the generated images with highest PSM” (PSM means perceptual similarity metrics, and it was calculated from early/middle/late layers of several image recognition deep neural networks [DNNs]), as illustrated in Fig 3A. This selection might lead readers or peer reviewers, particularly those not specialized in the brain decoding field, to overestimate the effectiveness of the methods and potentially lead to a distorted understanding of the actual reconstruction performance. Note that the Brain-Diffuser study [45] did not execute such procedures, and in their subsequent report [53], they updated the image presentation procedure more fairly as: “we generated five images with different stochastic noise and selected three images randomly.”
@@
Fig 3: Analysis of post-hoc selection in the StableDiffusionReconstruction method [44]. (A) Selection procedure. Five images were generated, and the one most closely resembling the test image was selected based on perceptual similarity metrics (PSM). (B) Images were generated from the CLIP text features translated from original brain activity and the VAE (vision) features from shuffled brain activity. Examples selected from five generations are shown with the test images. (C) Images were generated from the CLIP text and the VAE (vision) features, both translated from shuffled brain activity. Examples of selected images are shown with the text images. Examples are shown as in (B).
@@
To examine the impact of this post-hoc selection procedure, we conducted an experiment using random brain data. Instead of feeding the test brain data into trained translators, we shuffled the activities within each voxel of the NSD test set independently to create random brain data. Specifically, the brain data of the NSD test set is in a matrix shape, with rows representing stimulus samples and columns representing voxels. To generate the random brain data, we selected one column (voxel) of the matrix and randomly shuffled its values. This process was repeated for all voxels independently. Surprisingly, when the random brain data were input into the VAE feature translator, which contributes to producing initial images, plausible images were obtained by generating five images and selecting the best one (Fig 3B). Even more strikingly, when the random brain data were input into both VAE and CLIP text feature translators, we still obtained convincing results by simply generating images five times and conducting the selection mentioned above (Fig 3C). These observations are inexplicable because the artificially created brain data should completely lack any information related to the original visual stimuli.